1. A gas system with constant pressure has change in internal energy of 240 J. Heat was lost to the surroundings and 150 J work was done to the surroundings as well. How many calories of heat was lost?
1
Expert's answer
2022-06-30T13:41:04-0400


∆U — change in internal energy = 240 J ;

Q — heat added to the system = ?

W — work done by the system = 150 J


∆U = Q - W


Q = ∆U + W = 240 + 150 = 390 J =

= 93.21 calories


Answer: 93.21 calories